# Kac–Moody Symmetric Spaces: arbitary symmetrizable complex or almost split real type

### v3

@@ 3.3 Twin Buildings
The pair $(\mathcal{W},S)$ with generating set $S=\{r_{1},...,r_{n}\}$ is a Coxeter system. Note that any Weyl group of a Kac–Moody algebra forms a Coxeter system and the order of the elements are in relation with the entries of the generalized Cartan matrix, see [Mar18, Proposition 4.22].
Definition 3.8 (cp. Definition 5.1 from [AB08])
@@
In order to associate a building to a group, we need the notion of a $BN$-pair, which implies the Bruhat decomposition. An important decomposition in order to define the structure of a building on a group, see [AB08, Section 6.2.6].
@@
where $\tilde{w}$ is a representative of the preimage of $w$ in the extended Weyl group under the surjection $\widetilde{\mathcal{W}}\to\mathcal{W}$, see Equation 3. This explicit description allows us to define a decomposition, as we will see later.
@@
Definition 3.9 (cp. Definition 6.55 from [AB08])
@@
    • for all $w\in N/T$ and $s\in S$ is $sBw\subseteq BswB\cup BwB$, and
    • $sBs^{-1}$ is no subgroup of $B$ for all $s\in S$.
We denote by $T$ the intersection $N\cap B$ and by $\mathcal{W}$ the quotient $N/T$. That a group $G$ has a $BN$-pair $(B,N)$ implies the Bruhat decomposition
@@
see [AB08, Section 6.2.6] By this decomposition and using Equation 4 follows that there is the desired bijection
@@
Given two buildings $\Delta_{+}=(\mathcal{C}_{+},\delta_{+})$ and $\Delta_{-}=(\mathcal{C}_{-},\delta_{-})$ of type $(\mathcal{W},S)$, where $\mathcal{C}_{-}\cap\mathcal{C}_{+}=\emptyset$. Then call the triple
$\Delta_{\pm}\coloneqq(\mathcal{C}_{+},\mathcal{C}_{-},\delta^{\star}),$
@@
$b\delta^{\star}\colon(\mathcal{C}_{+}\times\mathcal{C}_{-})\cup(\mathcal{C}_{-}\times\mathcal{C}_{+})\to\mathcal{W}$
@@
  • $\delta^{\star}(C,D)=\delta^{\star}(D,C)^{-1}$.
  • If $\delta^{\star}(C,D)=w\in\mathcal{W}$ and if $\delta_{\varepsilon}(E,C)=s\in S$ for $E\in\mathcal{C}_{\varepsilon}$ and $\langle(sw)<l(w)$, then $\delta^{\star}(E,D)=sw$.
@@
a twin building of type $(\mathcal{W},S)$ with codistance function $\delta^{\star}$. For more details, see [AB08, Definition 5.133].
By [AB08, Definition 6.78] can a building constructed from a group $G$ if it has a twin $BN$-pair $(B_{+},B_{-},N)$, which is a $BN$-pair with respect to the subgroup $B_{+}$ and $B_{-}$. Similar to the case for one building, a twin $BN$-pair guaranties the Birkhoff decomposition (cf. [AB08, Proposition 6.81]):
@@
where $\varepsilon\in\{+,-\}$. Again, as in the case for one building, this decomposition allows to define a codistance function on the set of chambers $G/B_{\pm}$. Note that $B_{+},B_{-}$ and $N$ are subgroups of $G$ satisfying some properties, for details see [AB08, Definition 6.78].
@@
Since we are interested in Kac–Moody groups, the notion of an RGD-system is useful, see [AB08, Section 8.6.1] and [AB08, Definition 7.82] for details. An RGD system allows to construct a $BN$-pair and thus associate a building to a group.
@@
    For all $\alpha,\beta\in\Delta^{\mathrm{re}}$ with $\alpha\neq\beta$ such that the set $\{\alpha,\beta\}$ is prenilpotent (see [AB08, Definition 8.41]), follows that $[U_{\alpha},U_{\beta}]\leq U_{(\alpha,\beta)}$, where $(\alpha,\beta)\coloneqq[\alpha,\beta]\setminus\{\alpha,\beta\}$.
  3. RGD.3
    For every $s\in S$ there exists a function $m\colon U_{\alpha_{s}}\setminus\{e\}\to G$, such that for all $u\in U_{\alpha_{s}}\setminus\{e\}$ and $\alpha\in\Delta^{\mathrm{re}}$ follows that
@@
    For all $s\in S$ is $U_{-\alpha_{s}}$ no subgroup of $U_{+}$.
@@
    $G=T\langle U_{\alpha}\mid\alpha\in\Delta^{\mathrm{re}}\rangle$
  6. RGD.6
@@
see [AB08, Remark 7.87]. Further in [AB08, Chapter 8.8] it is shown, that $N_{G}(U_{\pm})=B_{\pm}$ and that $B_{+}\cap B_{-}=T$ ([AB08, Corollary 8.78]).
All together, it is shown in [AB08, Theorem 8.80] that an RGD system leads to an $BN$-pair with bijection
@@ 3.4 A Characteristic Subgroup
A very useful subgroup to study symmetric spaces which belongs to a Kac–Moody group is the centered subgroup, that is the subgroup only generated by the root groups.
@@
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$. Denote with $G^{\dagger}<G$ the subgroup of $G$ generated by the root subgroups of $G$, i.e.
@@
The group is called a centered subgroup.
Note, that from 3.10 it is clear, that $G^{\dagger}$ is a normal subgroup of $G$. From [AB08, Corollary 8.79] it follows that the torus of a Kac–Moody group can be described as the intersection of all normalizers of the root groups. We use this fact to define a torus for $G^{\dagger}$:
@@
Since the RGD axioms are manly using root groups, it follows that $(G^{\dagger},(U_{\alpha})_{\alpha\in\Delta^{\mathrm{re}}},T^{\dagger})$ forms an RGD system. Moreover, according to [AB08, Proposition 8.82], the kernel of the action of $G$ on the associated twin building $\Delta_{\pm}$ is contained in the torus $T$, so that
@@
Hence the type of the twin building associated to $G^{\dagger}$ is the same as the type of the twin building associated to $G$.
@@
By the RGD axioms one has $G=TG^{\dagger}$ (see [AB08, Chapter 8]). Moreover, $G^{\dagger}$ is a normal subgroup of $G$, since
@@
cp. [AB08, Corollary 8.79]. Furthermore $G/G^{\dagger}$ is abelian, therefore $[G,G]\subseteq G^{\dagger}$.
@@
$G_{i}=\langle U_{\alpha_{i}},U_{-\alpha_{i}}\rangle\cong\operatorname{SL}(2,k).$
Due to the fact that $\operatorname{SL}(2,k)$ is perfect for $k\in\{\mathbb{R},\mathbb{C}\}$, one deduces that $G^{\dagger}$ is perfect. Thus
@@
Since we are interested in a generalization of [FHHK20], it is useful to know that the centered subgroup $G^{\dagger}$ corresponds to the type of Kac-Moody group used by the authors in [FHHK20] to develop the theory of symmetric Kac–Moody spaces. In particular, the type of Kac–Moody groups used in [FHHK20] is called centered, which means that the group is generated only by its root subgroups, without an additional torus. In the language of root data, the group is generated by a simply connected root datum, cf. [Mar18, Example 7.11].
@@ 3.5 The Chevalley Involution
The first goal of this article is to construct a symmetric space with respect to a Kac–Moody group. For this purpose, it is useful and important to define an involution. Therefore let $G$ be a Kac–Moody group over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ of type $\mathbb{A}$. Then $\Theta\colon G\to G$ denotes a continuous involution which is given for $\mathbb{K}=\mathbb{C}$ by
@@
for $r\in\mathbb{C}$ and $\alpha\in\Delta^{\mathrm{re}}$. Since the complex conjugation acts trivially on the reals, it can be omitted in the case $\mathbb{K}=\mathbb{R}$. In [Cap09, §8.2] it is described how this involution is obtained by lifting the Chevalley involution of a Kac–Moody algebra $g(\mathbb{A})$ (cf. [Kac90, §1.3]) to a continuous involution on $G$. In particular, this means that $\Theta$ maps $U_{\alpha}$ to $U_{-\alpha}$ and stabilizes the torus $T$.
Denote by $K=G^{\Theta}$ the set of fixed points of $\Theta$ and define by $\tau\colon G\to G$, $g\mapsto g\Theta(g)^{-1}$ the twist map of $G$. In the following lemma, we define useful subgroups and analyze their properties. The statements are based on [FHHK20, Lemma 3.24] and [FHHK20, Lemma 3.26].
@@
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$, where $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$. Define the subgroups $A\coloneqq\tau(T)$ and $M\coloneqq K\cap T$. Then it follows that $T\cong M\times A$ and the following properties are satisfied
@@
In order to prove (A2), we first check, that for an element $t\in T$ holds
$\tau(t)=\begin{cases}t^{2}&\text{for }\mathbb{K}=\mathbb{R}\\ |t|^{2}&\text{for }\mathbb{K}=\mathbb{C}.\end{cases}$
This follows directly from the definition of the action of $\Theta$ on $T$, to be precise, for $t\in T$ applies $\Theta(t)=\overline{t}^{-1}$ for $\mathbb{K}=\mathbb{C}$ and $\Theta(t)=t^{-1}$ for $\mathbb{K}=\mathbb{R}$. Hence, it follows that $A=\tau(T)\cong(\mathbb{R}_{>0})^{2n-l}$, since the dimension of $T$ is $2n-l$.
Next, we want to check (A4). Let $g\in K\cap A\subseteq T$. Then by definition of $K$ and $\Theta$ it follows that
@@
Thus $g\overline{g}=e$ and from the fact that $A=\tau(T)\cong(\mathbb{R}_{>0})^{2n-l}$ it follows that $g=e$. In particular, the statement that $M\cap A=\{e\}$ is true follows then directly from $M\subset K$.
Finally, we prove the statement of (A5). Therefore recall, that by the definition of $B_{\pm}$ it follows that $\Theta(B_{\pm})=B_{\mp}$ and that $T=B_{+}\cap B_{-}$. Note, that this is based on the fact that $U_{+}\cap U_{-}=\{e\}$. Thus, it follows that $K\cap B_{\pm}=K\cap T$. Now let $g\in K\cap B_{\pm}$, then follows that $g\in T$. Using the decomposition $T\cong M\times A$ we can write $g=ma$ for $m\in M$ and $a\in A$., one calculates
$\tau(g)=\begin{cases}g^{2}=m^{2}a^{2}=a^{2}&\text{for }\mathbb{K}=\mathbb{R}\\ |g|^{2}=|m|^{2}|a|^{2}=|a|^{2}=a^{2}&\text{for }\mathbb{K}=\mathbb{C}.\end{cases}$
On the other hand, by the fact that $g\in K$ follows that $\tau(g)=e$. This leads to $a^{2}=e$ and since $a\in A\cong(\mathbb{R}_{>0})^{2n-l}$ it follows that $\tau(g)=e$ if and only if $a=e$.
The fact that $T\cong M\times A$ follows from the polar decomposition:
@@ 3.6 The Iwasawa Decomposition
In the theory of symmetric spaces, it is necessary to define flats because they are the most important principles for understanding actions and causal boundaries. To do so, one needs the topological Iwasawa decomposition, i.e. that the multiplication induces a homeomorphism
@@
First we will prove the algebraic statement that for a Kac–Moody group $G$ there is a decomposition $G=U_{\pm}AK$ which induces a bijection. After establishing the algebraic Iwasawa decomposition, we will use the same strategy as in the proof of [FHHK20, Theorem 3.31 (ii)] to prove the topological assertion for Kac–Moody groups over $\mathbb{K}$ of arbitrary type $\mathbb{A}$.
@@
$=U_{\pm}(MA)A^{\dagger}K^{\dagger}=U_{\pm}AK$
since $K\setminus K^{\dagger}$ can by only contained in $T$ and by 3.14 (A5) it is contained in $M$. The other order of the decomposition results from the symmetry arguments.
@@
Hence $a_{1}^{-1}u_{1}^{-1}u_{2}a_{2}\in K$ and $B_{\pm}\cap K=M\subseteq T$, see 3.14. It follows that
@@
(cp. [AB08, Lemma 7.62]) and from this, that
@@
It follows that $k_{1}=k_{2}$ and thus $m_{\pm}$ is injective. ∎
@@
Now for the topological version, the idea is to use the same strategy as in the proof of [FHHK20, Theorem 3.31 (ii)], to be precise we will construct the inverse functions and show that they are open. That we can use the same strategy as in [FHHK20] is based on the statement that every Kac–Moody group has a strong topological building, as shown by Grüning and the first-named author of this article in [HK23, Appendix B] for centered Kac–Moody groups. The very technical generalization for arbitrary split Kac–Moody groups can be found in [Voc23, Appendix C]. For an introduction to topological buildings see [HKM13, Section 3.1] and for strong topological buildings in particular see [HKM13, Definition 3.21].
@@
is trivial.
@@
First, note that $G$ leads to a non-spherical strong topological twin building, according to a result by Grüning and the first-named author of this article, see [HK23, Corollary B.8] (or [Voc23, Appendix C]). Furthermore, note that $T\cong A\times M$ as topological groups and that $U_{\mp}\times T\times U_{\pm}$ is open. In particular the map $U_{\mp}\times T\times U_{\pm}\to B_{\mp}B_{\pm}$ is a homeomorphism (cf. [AB08, Section 8.8] and [HK23, Corollary B.7]) and $B_{\mp}B_{\pm}$ is open (cf. [HKM13, Lemma 6.1]). Hence
@@
and any open subset $Y\subseteq X$ is open in $X$ if and only if for every $n\in\mathbb{N}$ $Y\cap X_{n}$ is open in $X_{n}$. Call the pair $(X,(X_{n})_{n\in\mathbb{N}})$ $k_{\omega}$-pair. For more details about $k_{\omega}$-spaces see [FT77] or [FHHK20, Definition 3.33 et seq.]. Here we recall two lemmas about $k_{\omega}$-spaces for the reader’s convenience that we use in order to derive the topological Iwasawa decomposition.
@@
Let $(X,(X_{n}))$ and $(Y,(Y_{n}))$ be $k_{\omega}$ pairs and let $f\colon X\to Y$ be a continuous bijection such that for all $n\in\mathbb{N}$ there exists a $m\in\mathbb{N}$ such that $Y_{n}\subset f(X_{m})$.
@@
Let $(X,(X_{n}))$ be a $k_{\omega}$ pair and let $\pi\colon\widetilde{X}\to X$ be a finite-sheeted covering. Then $(\widetilde{X},\pi^{-1}(X_{n}))$ is a $k_{\omega}$ pair.
@@
$G_{l}^{\pm}\coloneqq\bigcup_{w\in\mathcal{W},\,l(w)\leq l}B_{\pm}wB_{\pm},$
and denote with $\widetilde{D}_{l,\pm}$ the image of $G_{l}^{\pm}$ in $G/AU_{\pm}$ and with $D_{l,\pm}$ the image of $G_{l}^{\pm}$ in $G/B_{\pm}$. By [HKM13, Corollary 3.13] follows that
@@
is a $k_{\omega}$-pair. Note that the preimage of $D_{l,\pm}$ under the fibration $p_{\pm}$ is exactly $\widetilde{D}_{l,\pm}$ and that the fibration by 3.16 is trivial, so we can apply 3.18, which works here, since the statement is just a topological assertion about $k_{\omega}$ pairs, so we observe that
@@
To conclude that $(K,(K_{l}^{\pm}))$ is a $k_{\omega}$ pair, we must show that the subsets $(K_{l}^{\pm})$ are compact. Therefore, recall that $G$ can be written as $KB_{\pm}$ using 3.15, and hence $K$ acts chamber-transitive on the to $G$ associated twin building $\Delta_{\pm}$. By the construction of $G_{l}^{\pm}$ and $K_{l}^{\pm}$ it follows that
$G_{l}^{\pm}=K_{l}^{\pm}(G_{l}^{\pm}\cap B_{\pm}).$

### v4

@@ 3.3 Twin Buildings
The pair $(\mathcal{W},S)$ with generating set $S=\{r_{1},...,r_{n}\}$ is a Coxeter system. Note that any Weyl group of a Kac–Moody algebra forms a Coxeter system and the order of the elements are in relation with the entries of the generalized Cartan matrix, see [Mar18, Proposition 4.22]. In detail, any finite Coxeter group is called a Weyl group.
Definition 3.8 (cf. Definition 5.1 from [AB08])
@@
In order to associate a building to a group, we need the notion of a $BN$-pair, which implies the Bruhat decomposition, for details see [AB08, Section 6.2.6]. It should also be noted that in this context of $BN$-pairs the abstract Coxeter group as we defined above becomes a Weyl group in the sense of a Kac–Moody group (for the definition see the end of Section 3.1).
@@
where $\tilde{w}$ is a representative of the preimage of $w$ in the extended Weyl group under the surjection $\widetilde{\mathcal{W}}\to\mathcal{W}$, see Equation 3. Note that $\mathcal{W}$ stands for a Weyl group in this context. This explicit description allows us to define a decomposition, as we will see later.
@@
Definition 3.9 (cf. Definition 6.55 from [AB08])
@@
    • for all $w\in N/T$ and $s\in S$ we have $sBw\subseteq BswB\cup BwB$, and
    • $sBs^{-1}$ is not a subgroup of $B$ for all $s\in S$.
We denote by $T$ the intersection $N\cap B$ and by $\mathcal{W}$ the quotient $N/T$. If a group $G$ has a $BN$-pair $(B,N)$, then the Bruhat decomposition holds:
@@
see [AB08, Section 6.2.6]. This decomposition combined with Equation 4 yield the desired bijection
@@
Given two buildings $\Delta_{+}=(\mathcal{C}_{+},\delta_{+})$ and $\Delta_{-}=(\mathcal{C}_{-},\delta_{-})$ of type $(\mathcal{W},S)$, where $\mathcal{C}_{-}\cap\mathcal{C}_{+}=\emptyset$. A twin building of type $(\mathcal{W},S)$ with codistance function $\delta^{\star}$ is a triple
$\Delta_{\pm}\coloneqq(\mathcal{C}_{+},\mathcal{C}_{-},\delta^{\star})$
@@
$\delta^{\star}\colon(\mathcal{C}_{+}\times\mathcal{C}_{-})\cup(\mathcal{C}_{-}\times\mathcal{C}_{+})\to\mathcal{W}$
@@
  • $\delta^{\star}(C,D)=\delta^{\star}(D,C)^{-1}$;
  • if $\delta^{\star}(C,D)=w\in\mathcal{W}$ and if $\delta_{\varepsilon}(E,C)=s\in S$ for $E\in\mathcal{C}_{\varepsilon}$ and $l(sw)<l(w)$, then $\delta^{\star}(E,D)=sw$;
@@
For more details, see [AB08, Definition 5.133].
By [AB08, Definition 6.78] a building can be constructed from a group $G$ if it has a twin $BN$-pair $(B_{+},B_{-},N)$, which is a $BN$-pair with respect to the subgroup $B_{+}$ and $B_{-}$. Similar to the case for a single building, a twin $BN$-pair guarantees the Birkhoff decomposition (cf. [AB08, Proposition 6.81]):
@@
where $\varepsilon\in\{+,-\}$. Again, as in the case of a single building, this decomposition allows to define a codistance function on the set of chambers $G/B_{\pm}$. Note that $B_{+},B_{-}$ and $N$ are subgroups of $G$ satisfying some properties, for details see [AB08, Definition 6.78].
@@
The notion of an RGD system is very useful for studying Kac–Moody group. For instance, an RGD system allows one to construct a $BN$-pair and, thus, associate a building to a group. Below we recall the definition of an RGD system, see also [AB08, Section 8.6.1] and [AB08, Definition 7.82]. In this article we will only need a few properties of an RGD system, which is why we will not go into the technical details here.
@@
    For all $\alpha,\beta\in\Delta^{\mathrm{re}}$ with $\alpha\neq\beta$ such that the set $\{\alpha,\beta\}$ is prenilpotent, we have $[U_{\alpha},U_{\beta}]\leq U_{(\alpha,\beta)}$, where $(\alpha,\beta)\coloneqq[\alpha,\beta]\setminus\{\alpha,\beta\}$. For details on the notation see [AB08, Definition 8.41].
@@
    For every $s\in S$ there exists a function $m\colon U_{\alpha_{s}}\setminus\{e\}\to G$, such that for all $u\in U_{\alpha_{s}}\setminus\{e\}$ and $\alpha\in\Delta^{\mathrm{re}}$ we have
@@
    $G=T\langle U_{\alpha}\mid\alpha\in\Delta^{\mathrm{re}}\rangle$
@@
see [AB08, Remark 7.87]. Further in [AB08, Chapter 8.8] it is shown, that $N_{G}(U_{\pm})=B_{\pm}$ and that $B_{+}\cap B_{-}=T$ ([AB08, Corollary 8.78]). Altogether, by [AB08, Theorem 8.80] an RGD system leads to a $BN$-pair with bijection
@@ 3.4 A Characteristic Subgroup
A very useful subgroup for studying symmetric spaces associated to Kac–Moody group is the centered subgroup, that is the subgroup only generated by the root groups.
@@
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$. Denote by $G^{\dagger}<G$ the subgroup of $G$ generated by the root subgroups of $G$, i.e.
@@
This subgroup is called the centered subgroup.
We can deduce that $G^{\dagger}$ is a normal subgroup of $G$: By RGD.4 we can write
$G=TG^{\dagger}.$
To deduce, that $G^{\dagger}$ is normal, consider an element $g\in G$ which we can also write as $g=t\widetilde{g}$. Now compute
$gG^{\dagger}g^{-1}=t\widetilde{g}G^{\dagger}\widetilde{g}^{-1}t^{-1}.$
By definition of $G^{\dagger}$ and by RGD.5 follows, that $gG^{\dagger}g^{-1}\subseteq G^{\dagger}$. From [AB08, Corollary 8.79] follows that the torus of a Kac–Moody group can be described as the intersection of all normalizers of the root groups. We use this fact to define a torus for $G^{\dagger}$:
@@
Since the RGD axioms are manly using root groups, we can deduce that $(G^{\dagger},(U_{\alpha})_{\alpha\in\Delta^{\mathrm{re}}},T^{\dagger})$ forms an RGD system. Moreover, according to [AB08, Proposition 8.82], the kernel of the action of $G$ on the associated twin building $\Delta_{\pm}$ is contained in the torus $T$, so that
@@
Note that we use the same arguments here as in a similar statement which can be found under the definition of an RGD system (see 3.10). By [AB08, Theorem 8.80], an RGD system leads to a $BN$-pair with this bijection.
We conclude this paragraph with the observation that the type of the twin building belonging to $G^{\dagger}$ is the same as the type of the twin building belonging to $G$.
@@
By RGD.4 one has $G=TG^{\dagger}$ (see [AB08, Chapter 8]). Moreover, $G^{\dagger}$ is a normal subgroup of $G$, since
@@
cf. [AB08, Corollary 8.79]. Since the natural map $T\mapsto G/G^{\dagger}$ is surjective, the quotient $G/G^{\dagger}$ is abelian and thus $[G,G]\subseteq G^{\dagger}$.
@@
$G_{i}=\langle U_{\alpha_{i}},U_{-\alpha_{i}}\rangle\cong\operatorname{SL}_{2}(\mathbb{K}).$
Due to the fact that $\operatorname{SL}_{2}(\mathbb{K})$ is perfect for $k\in\{\mathbb{R},\mathbb{C}\}$, one deduces that $G^{\dagger}$ is perfect. Thus
@@
Note that this centered subgroup $G^{\dagger}$ is the Kac-Moody group used by the authors in [FHHK20] to develop their theory of Kac–Moody symmetric spaces.
@@ 3.5 The Chevalley Involution
The first goal of this article is to construct a symmetric space with respect to a Kac–Moody group. For this purpose, it is useful and important to define an involution. Therefore, let $G$ be a Kac–Moody group over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ of type $\mathbb{A}$. Then $\Theta\colon G\to G$ denotes a continuous group involution which is given for $\mathbb{K}=\mathbb{C}$ by
@@
for $r\in\mathbb{C}$ and $\alpha\in\Delta^{\mathrm{re}}$. From this definition one can easily derive the following
$\Theta(t)=\overline{t}^{-1}$
for all $t\in T$.
The map $\Theta$ is continuous with respect to the Kac–Peterson topology on $G$. Since the complex conjugation acts trivially on the reals, it can be omitted in the case $\mathbb{K}=\mathbb{R}$. By [Cap09, §8.2] this involution can be obtained by lifting the Chevalley involution of a Kac–Moody algebra $g(\mathbb{A})$ (cf. [Kac90, §1.3]) to a continuous involution on $G$. In particular, this means that $\Theta$ maps $U_{\alpha}$ to $U_{-\alpha}$ and stabilizes the torus $T$.
Denote by $K=G^{\Theta}$ the set of fixed points of $\Theta$ and define by $\tau\colon G\to G$, $g\mapsto g\Theta(g)^{-1}$ the twist map of $G$. In the case that $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$, it can be derived using the definition of $\Theta$ that
$\tau(t)=\begin{cases}t^{2}&\text{for }\mathbb{K}=\mathbb{R}\\ |t|^{2}&\text{for }\mathbb{K}=\mathbb{C},\end{cases}$
for all $t\in T$, or simply $\tau(t)=|t|^{2}$.
In the following lemma, we define useful subgroups and analyze their properties. The statements are based on [FHHK20, Lemma 3.24] and [FHHK20, Lemma 3.26].
@@
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$, where $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$. Define the subgroups $A\coloneqq\tau(T)$ and $M\coloneqq K\cap T$. Then we have $T\cong M\times A$ and the following hold:
@@
The statement of (A2) follows directly from the definition of the maps, see the paragraph above. Concretely, this shows that $A=\tau(T)\cong(\mathbb{R}_{>0})^{2n-l}$, because the dimension of $T$ is $2n-l$.
Next, we want to check (A3). Let $g\in K\cap A\subseteq T$. Then by definition of $K$ and $\Theta$ follows that
@@
Thus $g\overline{g}=e$ and from the fact that $A=\tau(T)\cong(\mathbb{R}_{>0})^{2n-l}$ one concludes that $g=e$. In particular, the statement that $M\cap A=\{e\}$ is true follows then directly from $M\subset K$.
The statement of (A4) follows from (A3) and the polar decomposition
@@
Finally, we prove the statement of (A5). Therefore recall, that by the definition of $B_{\pm}$ one has $\Theta(B_{\pm})=B_{\mp}$ and that $T=B_{+}\cap B_{-}$. Note, that this is based on the fact that $U_{+}\cap U_{-}=\{e\}$. Thus we have $K\cap B_{\pm}=K\cap T$ = M, by definition. ∎∎
@@ 3.6 The Iwasawa Decomposition
In the theory of symmetric spaces, it is necessary to define flats because they are the most important objects for understanding actions and causal boundaries. To do so, one needs the topological Iwasawa decomposition, i.e. that the multiplication induces a homeomorphism
@@
First we will prove the algebraic statement that for a Kac–Moody group $G$ there is a decomposition $G=U_{\pm}AK$. After establishing the algebraic Iwasawa decomposition, we will use the same strategy as in the proof of [FHHK20, Theorem 3.31 (ii)] to prove the topological assertion for Kac–Moody groups over $\mathbb{K}$ of arbitrary type $\mathbb{A}$.
@@
$=U_{\pm}(MA)A^{\dagger}K^{\dagger}$
$=U_{\pm}MAK^{\dagger}$
$=U_{\pm}AMK^{\dagger}\subset U_{\pm}AK$
since $M,K^{\dagger}\subset K$ (see 3.14 (A5)). The other order of the decomposition results from the symmetry arguments.
@@
Hence $a_{1}^{-1}u_{1}^{-1}u_{2}a_{2}\in K$ and $B_{\pm}\cap K=M\subseteq T$, see 3.14. And as a result
@@
(cf. [AB08, Lemma 7.62]) and from this, that
@@
Therefore $k_{1}=k_{2}$ and thus $m_{\pm}$ is injective. ∎
@@
Now for the topological version, the idea is to use the same strategy as in the proof of [FHHK20, Theorem 3.31 (ii)], to be precise we will construct the inverse functions and show that they are open. The fact that we can use the same strategy as in [FHHK20] is based on the statement that every Kac–Moody group has a strong topological building, as shown by Grüning and the first-named author of this article in [HK23, Appendix B] for centered Kac–Moody groups. The very technical generalization for arbitrary split Kac–Moody groups can be found in [Voc23, Appendix C]. For an introduction to topological buildings see [HKM13, Section 3.1] and for strong topological buildings in particular see [HKM13, Definition 3.21].
@@
is trivial, i.e., a topological direct product of the fiber and the base space.
@@
First, note that $G$ leads to a non-spherical strong topological twin building, according to a result by Grüning and the first-named author of this article, see [HK23, Corollary B.8] (or [Voc23, Appendix C]). Furthermore, note that $T\cong A\times M$ as topological groups and that $U_{\mp}\times T\times U_{\pm}$ is open. In particular the map $U_{\mp}\times T\times U_{\pm}\to B_{\mp}B_{\pm}$ is a homeomorphism (cf. [HK23, Corollary B.7]) and $B_{\mp}B_{\pm}$ is open (cf. [HKM13, Lemma 6.1]). Hence
@@
and any open subset $Y\subseteq X$ is open in $X$ if and only if for every $n\in\mathbb{N}$ $Y\cap X_{n}$ is open in $X_{n}$. Call the pair $(X,(X_{n})_{n\in\mathbb{N}})$ $k_{\omega}$-pair. For more details about $k_{\omega}$-spaces see [FT77] or [FHHK20, Definition 3.33 et seq.]. Note that a Kac–Moody group over $\mathbb{K}$ equipped with the Kac–Peterson topology is a $k_{\omega}$-group, for a justification and details see [HKM13, Proposition 7.10] and [HK23, Definition and Remark 2.1].
Here we recall two lemmas about $k_{\omega}$-spaces for the reader’s convenience that we use in order to derive the topological Iwasawa decomposition.
@@
Let $(X,(X_{n}))$ and $(Y,(Y_{n}))$ be $k_{\omega}$-pairs and let $f\colon X\to Y$ be a continuous bijection such that for all $n\in\mathbb{N}$ there exists a $m\in\mathbb{N}$ such that $Y_{n}\subset f(X_{m})$.
@@
Let $(X,(X_{n}))$ be a $k_{\omega}$-pair and let $\pi\colon\widetilde{X}\to X$ be a finite-sheeted covering. Then $(\widetilde{X},\pi^{-1}(X_{n}))$ is a $k_{\omega}$-pair.
@@
$G_{l}^{\pm}\coloneqq\bigcup_{w\in\mathcal{W},\,l(w)\leq l}B_{\mp}wB_{\pm},$
and denote with $\widetilde{D}_{l,\pm}$ the image of $G_{l}^{\pm}$ in $G/AU_{\pm}$ and with $D_{l,\pm}$ the image of $G_{l}^{\pm}$ in $G/B_{\pm}$. By [HKM13, Corollary 3.13] we have
@@
is a $k_{\omega}$-pair. Note that the preimage of $D_{l,\pm}$ under the fibration $p_{\pm}$ is exactly $\widetilde{D}_{l,\pm}$ and that the fibration by 3.16 is trivial, so we can apply 3.18, which works here, since the statement is just a topological assertion about $k_{\omega}$-pairs, so we observe that
@@
To conclude that $(K,(K_{l}^{\pm}))$ is a $k_{\omega}$-pair, we must show that the subsets $(K_{l}^{\pm})$ are compact. To this end, recall that $G$ can be written as $KB_{\pm}$ using 3.15, and hence $K$ acts chamber transitively on the to $G$ associated twin building $\Delta_{\pm}$. In particular
$\Delta_{\pm}\quad\quad\text{and}\quad\quad K/(K\cap B_{\pm})=K/(K\cap T)$
are isomorphic as $K$-sets. Since $G_{l}^{\pm}$ is a union of $B_{\pm}$-double cosets and, moreover, is invariant under the Cartan-Chevalley involution, we have
$G_{l}^{\pm}=K_{l}^{\pm}B_{\pm}.$